OpenStudy (anonymous):

can anybody help me please the steps with these (6x+7)(3x+2)

OpenStudy (anonymous):

you have to do four multiplications \[(6x+7)(3x+2)=6x\times 3x+6x\times 2+7\times 3x+7\times 2\]

OpenStudy (amistre64):

just mulitply it out 6x +7 3x +2 -------------- 12x+14 18x^2 +21x --------------- 18x^2 +33x +14

OpenStudy (anonymous):

when you get done you will have \[18x^2+33x+14\]

OpenStudy (anonymous):

you can do it in one line like i did or multiply like you would two "two digit" numbers as amisre did

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

i wonder where is the 33 come from?

OpenStudy (anonymous):

oh here

OpenStudy (anonymous):

\[6x\times 2 + 7\times 3x=12x+21x=33x\]

OpenStudy (anonymous):

is the steps suppose to do these (6x+7)(3x+2) 12x+21x=33x 18x^2+21x =18x^2+33x+14